The width of a retangle is 15 feet less than its length, and the perimeter is 210 feet. Can you show me how to find the width and the length.
Answer by checkley71(8403)

W=L-15
P=2L+2W
210=2L+2(L-15)
210=2L+2L-30
4L=210+30
4L=240
L=240/4
L=60 THEN
W=60-15
W=45
PROOF
210=2*60+2*45
210=120+90
210=210